OpenStudy (anonymous):

Find the integral e^(3/x)/x^2

OpenStudy (tkhunny):

Have you considered a substitution?

OpenStudy (anonymous):

im really confused with u substitution

OpenStudy (anonymous):

im thinking that in this question: u=x^2 and du=2x dx right?

OpenStudy (anonymous):

how do you know which to substitute?

OpenStudy (tkhunny):

Try u = 1/x. This leads to du = -1/x^2 dx and vastly simplifies things. You must think of some significant part of the argument that will be the derivative of your substitution. In this case, I noticed that we have (1/x) in that exponent and the derivative of 1/x is -1/x^2 which just happens to match the denominator. Perfect!

OpenStudy (anonymous):

where are you gettin 1/x though?

OpenStudy (anonymous):

you mean 3/x?

OpenStudy (tkhunny):

No. 3/x = 3(1/x) You can use u 3/x if you like, you'll just get more constants to fix. If the problem were \(3\cdot\dfrac{e^{3/x}}{x^{2}}\), I would have picked u = 3/x. Note: It won't always be obvious. You have to think about it and try something. You will gain experience as you go along. Note: There may be more than one way to proceed. You do not have to find THE WAY. Find a way that works. Various constants don't matter. You can fix those on the fly.

OpenStudy (tkhunny):

* u = 3/x if you like

OpenStudy (anonymous):

if u=3/x then du=-3x^-2

OpenStudy (anonymous):

right?

OpenStudy (tkhunny):

\(du = -\dfrac{3}{x^{2}}dx\)

OpenStudy (anonymous):

so then\[\int\limits\frac{ e^u }{ x^2 } dx?\]

OpenStudy (tkhunny):

Well, sort of. You must do the rest. Substitute for \(\dfrac{dx}{x^{2}}\).

OpenStudy (tkhunny):

Given \(du = -\dfrac{3}{x^{2}}dx\), we have \(\dfrac{dx}{x^{2}} = \dfrac{du}{-3}\)

OpenStudy (anonymous):

and you take the integral of that? sorry im just really confused

OpenStudy (tkhunny):

Substitute ALL the pieces. The point is to make an integral that you can handle. We have: \(\int\dfrac{e^{3/x}}{x^{2}}\;dx\) We decide \(u = \dfrac{3}{x}\) We calculate: \(du = -\dfrac{3}{x^{2}}dx\) We rebuild the integral so the substitutionS will work perfectly \(\dfrac{1}{-3}\cdot\int\dfrac{e^{3/x}}{x^{2}}\cdot \left(-3\right)\;dx\) We substitute EVERTHING: \(\dfrac{1}{-3}\cdot\int\e^{u}\;du\) We solve the simpler integral: \(\dfrac{1}{-3}e^{u} + C\) We put things back: \(\dfrac{1}{-3}e^{3/x} + C\) Done.

OpenStudy (anonymous):

sorry to keep asking...but im still wondering how you got 1/-3? in the front of the integral

OpenStudy (tkhunny):

I just put it on the outside because I needed one on the inside. \(\int Stuff\;dx = \dfrac{1}{a}\int a\cdot Stuff\;dx\) I could have written: \(\int \dfrac{a}{a}\cdot Stuff \;dx\) You should know this principle: \(\int a\cdot f(x)\;dx = a\cdot\int f(x)\;dx\)

OpenStudy (anonymous):

ok that makes much more sense

OpenStudy (anonymous):

thanks for all your help and patience tkhunny :)

OpenStudy (tkhunny):

I needed WAY more partience for the connection problems than for you. Way to hang in there. Good work.

OpenStudy (anonymous):

thanks!